OpenStudy (anonymous):

determine graphically the x-value when y= -4 in the given equation. -x+2y= -6 how would this look on a graph when the given equation intersects the line y= -4

OpenStudy (anonymous):

-x+2y=-6 is a line

OpenStudy (anonymous):

can u tell me how it would look on a graph

OpenStudy (anonymous):

2y=-6+x y=x/2-3